OpenStudy (anonymous):

Use elimination to solve the system of equations. x+2y+3z=9 x+3y+2z=5 x+4y-z=-5

OpenStudy (e.mccormick):

Know what elimination means?

OpenStudy (e.mccormick):

here is an example: a + 2b = 5 -a + b = 1 If I add these, a gets eliminated: a + (-a) + 2b + (b) = 5 + (1) 0a +3b = 6 3b = 6 3b/3 = 6/3 b=2 Then you can can substitute that into either one to find a, but the elimination part allowed me to find b. A similar thing can be done to turn 3 equations in 3 unknowns into 2 in 2 unknowns, then those into 1 with 1 unknowns.